# Robust mass lumping and outlier removal strategies in isogeometric analysis

### v1

@@ 3.2 Block mass lumping
In general, for nontrivial geometries and coefficients, the mass matrix cannot be expressed as a Kronecker product and the techniques described in [22, 16, 12] are not straightforwardly applicable. In [32], the authors suggested computing the nearest Kronecker product approximation and substituting it to the consistent mass. However, the loss of accuracy induced by the approximation depends on the singular value decay of a reordered matrix and is independent of the discretization parameters. Not only may the approximation be rather crude but combining it with mass lumping also does not give a theoretical guarantee of improving the CFL, although it was observed in all practically relevant cases. Both issues will be addressed in this section. The key observation is that, although the mass matrix generally cannot be expressed as a Kronecker product, it still inherits some favorable structure from the tensor product basis functions which we will exploit for designing algebraic mass lumping techniques. We will first describe this structure in detail for the single-patch case and then propose a block generalization of the methods presented in [32]. From there, the multipatch case naturally follows.
@@
This hierarchical notion of bandedness is related to the standard notion of bandedness through the bandwidths and block sizes. In the rest of the paper we will denote $r_{k}=\prod_{j=k+1}^{d}n_{j}$ the size of the matrices on the $k$th hierarchical level (with $k<d$). On the finest level, the “blocks” reduce to scalars and we set $r_{d}=1$.
@@
Definitions 3.3 and 3.4 only provide a description of the sparsity. In order to build a working framework, they must be complemented with Definitions 3.6 and 3.7 that describe both symmetry and spectral properties.
@@
$\mathcal{S}_{(n_{1},n_{2},\dots,n_{d})}=\{\mathcal{B}\in\mathcal{S}_{n}\colon B_{i,j}\in\mathcal{S}_{(n_{2},\dots,n_{d})}\}\quad\text{and}\quad\mathcal{S}_{(n_{1},n_{2},\dots,n_{d})}^{+}=\{\mathcal{B}\in\mathcal{S}_{n}^{+}\colon B_{i,j}\in\mathcal{S}_{(n_{2},\dots,n_{d})}\}$
@@
$\mathcal{S}_{(n_{1},n_{2},\dots,n_{d})}\subseteq\mathcal{S}_{(n_{1},n_{2},\dots,n_{d-1},r_{d-1})}\subseteq\mathcal{S}_{(n_{1},n_{2},\dots,n_{d-2},r_{d-2})}\subseteq\dots\subseteq\mathcal{S}_{(n_{1},r_{1})}\subseteq\mathcal{S}_{n},$
$\mathcal{S}_{(n_{1},n_{2},\dots,n_{d})}^{+}\subseteq\mathcal{S}_{(n_{1},n_{2},\dots,n_{d-1},r_{d-1})}^{+}\subseteq\mathcal{S}_{(n_{1},n_{2},\dots,n_{d-2},r_{d-2})}^{+}\subseteq\dots\subseteq\mathcal{S}_{(n_{1},r_{1})}^{+}\subseteq\mathcal{S}_{n}^{+}.$
@@
We prove the inclusions from left to right. Let $\mathcal{B}\in\mathcal{S}_{(n_{1},n_{2},\dots,n_{d})}$. By definition, on the $(d-1)$th hierarchical level the matrices are in $\mathcal{S}_{n_{d}}=\mathcal{S}_{r_{d-1}}$ and the first inclusion trivially follows. On the $(d-2)$th level the matrices are in $\mathcal{S}_{(n_{d-1},r_{d-1})}\subseteq\mathcal{S}_{r_{d-2}}$. Thus, $\mathcal{B}\in\mathcal{S}_{(n_{1},n_{2},\dots,n_{d-2},r_{d-2})}$. We then repeatedly apply the same argument by moving up the hierarchical structure and realizing that at level $k$ (with $21\leq k\leq d-$ 2) the matrices are in $\mathcal{S}_{(n_{k+1},r_{k+1})}\subseteq\mathcal{S}_{r_{k}}$. The proof of the second statement is completely analogous. ∎
@@
$\mathbf{x}^{T}(\mathcal{L}(\mathcal{B})-\mathcal{B})\mathbf{x}$ $=\sum_{i=1}^{n_{1}}\mathbf{x}_{i}^{T}\left(\sum_{j=1}^{n_{1}}B_{i,j}\right)\mathbf{x}_{i}-\sum_{i,j=1}^{n_{1}}\mathbf{x}_{i}^{T}B_{i,j}\mathbf{x}_{j}$
@@
$\lambda_{k}(\mathcal{A},\mathcal{L}(\mathcal{B}))=\lambda_{k}(\mathcal{A},\mathcal{P}_{1})\leq\lambda_{k}(\mathcal{A},\mathcal{P}_{2})\leq\dots\leq\lambda_{k}(\mathcal{A},\mathcal{P}_{n_{1}-1})\leq\lambda_{k}(\mathcal{A},\mathcal{P}_{n_{1}})=\lambda_{k}(\mathcal{A},\mathcal{B})\qquad 1\leq k\leq n.$
Clearly, block diagonal matrices such as $\mathcal{P}_{1}$ are very appealing given that linear systems can be solved in parallel for each block. The block tridiagonal case can still be treated efficiently by a block forward elimination and backward substitution algorithm; see [45, Section 4.5.1] for the details.
@@ 3.3 Hierarchical mass lumping
Figure 3.2, for example, shows the sparsity pattern of a matrix $\mathcal{B}\in\mathcal{S}_{(n_{1},n_{2},n_{3})}^{+}$ together with its hierarchical lumped mass matrices $\mathcal{H}_{k}$ for $k=1,2,3$. By construction, hierarchical mass lumping reduces the bandwidth down the hierarchical structure: if $\mathcal{B}$ is $d$-level banded with bandwidths $(b_{1},b_{2},\dots,b_{d})$, then $\mathcal{H}_{k}$ is $d$-level banded with bandwidths $(0,\dots,0,b_{k+1},\dots,b_{d})$. Similarly to (3.5), hierarchical lumped matrices also satisfy an order relation.
@@ 3.4 Multipatch mass lumping
First recall that for any symmetric matrices $A,B\in\mathbb{R}^{n\times n}$ and any $V\in\mathbb{R}^{n\times m}$, if $A\succeq B$, then $V^{T}AV\succeq V^{T}BV$ [46, Theorem 7.7.2(a)]. The result then immediately follows since for any $1\leq r\leq N_{p}$ and any index $1\leq i<n_{1}$,
@@
The statement then follows from an inductive application of the previous inequality.
@@
The purpose of mass lumping is first and foremost to reduce the block bandedness of the mass matrix and guarantee a CFL condition that cannot be worse than the original one. However, generally speaking, mass lumping does not significantly improve the CFL while it might undermine the accuracy. Multipatch problems are not exempt and the issue already originates on the single-patch level, before local matrices are merged into the global one. This merging process is identical to the assembly procedure of classical finite element methods, which is not surprising given the analogy between patches and elements. Thus, the proof of the following lemma is well-known (see e.g. [47, 48]), but is repeated for the sake of completeness and notational consistency.
@@
Proof.
We first note that the assembly of global multipatch matrices can be written out more compactly as
$\mathcal{A}=\sum_{r=1}^{N_{p}}R_{r}^{T}\mathcal{A}_{r}R_{r}=\mathsf{R}^{T}\mathsf{A}\mathsf{R},\qquad\mathcal{B}=\sum_{i=1}^{N_{p}}R_{r}^{T}\mathcal{B}_{r}R_{r}=\mathsf{R}^{T}\mathsf{B}\mathsf{R}$
where $\mathsf{R}^{T}=[R_{1}^{T},\dots,R_{N_{p}}^{T}]$, $\mathsf{A}=\operatorname{\operatorname{diag}}(\mathcal{A}_{1},\dots,\mathcal{A}_{N_{p}})$ and $\mathsf{B}=\operatorname{\operatorname{diag}}(\mathcal{B}_{1},\dots,\mathcal{B}_{N_{p}})$. Consequently, by the Courant-Fischer theorem [46, Theorem 4.2.6],
$\lambda_{\min}(\mathcal{A},\mathcal{B})=\min_{\mathbf{x}\neq\mathbf{0}}\frac{\mathbf{x}^{T}\mathsf{R}^{T}\mathsf{A}\mathsf{R}\mathbf{x}}{\mathbf{x}^{T}\mathsf{R}^{T}\mathsf{B}\mathsf{R}\mathbf{x}}=\min_{\begin{subarray}{c}\mathbf{y}\in\mathcal{V}\\ \mathbf{y}\neq\mathbf{0}\end{subarray}}\frac{\mathbf{y}^{T}\mathsf{A}\mathbf{y}}{\mathbf{y}^{T}\mathsf{B}\mathbf{y}}\geq\min_{\mathbf{y}\neq\mathbf{0}}\frac{\mathbf{y}^{T}\mathsf{A}\mathbf{y}}{\mathbf{y}^{T}\mathsf{B}\mathbf{y}}=\min_{r}\lambda_{\min}(\mathcal{A}_{r},\mathcal{B}_{r}).$
where $\mathcal{V}$ is the space spanned by the columns of $\mathsf{R}$. Similarly,
$\lambda_{\max}(\mathcal{A},\mathcal{B})=\max_{\mathbf{x}\neq\mathbf{0}}\frac{\mathbf{x}^{T}\mathsf{R}^{T}\mathsf{A}\mathsf{R}\mathbf{x}}{\mathbf{x}^{T}\mathsf{R}^{T}\mathsf{B}\mathsf{R}\mathbf{x}}=\max_{\begin{subarray}{c}\mathbf{y}\in\mathcal{V}\\ \mathbf{y}\neq\mathbf{0}\end{subarray}}\frac{\mathbf{y}^{T}\mathsf{A}\mathbf{y}}{\mathbf{y}^{T}\mathsf{B}\mathbf{y}}\leq\max_{\mathbf{y}\neq\mathbf{0}}\frac{\mathbf{y}^{T}\mathsf{A}\mathbf{y}}{\mathbf{y}^{T}\mathsf{B}\mathbf{y}}=\max_{r}\lambda_{\max}(\mathcal{A}_{r},\mathcal{B}_{r}).$
The upper bound of Lemma 3.22 may be quite tight and is an incentive for acting on the single-patch matrices before assembling the global ones.
@@ 3.5 Solving linear systems with the lumped mass matrix
In the single-patch case, linear systems with the lumped mass matrices are conveniently solved using sparse Cholesky factorizations; i.e. $P=LL^{T}$, where $L$ is a lower triangular matrix. In isogeometric analysis, the fill-in of the Cholesky factor is well described by the envelope of the matrix [49].
Definition 3.23.
The envelope of a matrix $B\in\mathbb{R}^{n\times n}$ is defined as
$\operatorname{\operatorname{env}}(B)=\{(i,j)\colon 1\leq i\leq n,\ J_{i}(B)\leq j<i\},\quad J_{i}(B)=\min\{j\colon 1\leq j\leq i,\ b_{ij}\neq 0\}.$
In words, $J_{i}(B)$ is simply the index of the first nonzero entry in the $i$th row of the lower triangular part of $B$. The envelopes of the consistent mass and lumped mass matrices of Figure 3.2 are shown in Figure 3.3. It is well-known that any fill-in of the Cholesky factor can only occur within the envelope [50]. For this reason, many techniques for minimizing the fill-in are based on minimizing the envelope of a permuted matrix. The mass lumping techniques discussed in this work reduce the bandwidth and envelope of the consistent mass, thereby significantly accelerating sparse direct solvers. Indeed, a direct application of Lemma 3.5 shows that the bandwidth of $\mathcal{H}_{k}$ is
@@
(a) $\operatorname{\operatorname{env}}(\mathcal{M})$
(b) $\operatorname{\operatorname{env}}(\mathcal{H}_{1})$
(c) $\operatorname{\operatorname{env}}(\mathcal{H}_{2})$
(d) $\operatorname{\operatorname{env}}(\mathcal{H}_{3})$
Figure 3.3: Envelope of the matrices in Figure 3.2
@@
where $S=X-C^{T}D^{-1}C$ is the Schur complement of $D$ in $\mathcal{P}$ and $\tilde{\mathbf{g}}=\mathbf{g}-C^{T}D^{-1}\mathbf{f}$. Contrary to the consistent mass matrix (which features a similar structure), the Schur complement of the lumped mass matrix can be formed explicitly and cheaply owing to its sparsity and simple block diagonal structure. Once the Schur complement is formed, which is done once and for all, (3.6) can be solved by backward substitution.
@@ 4.1 Deflation techniques
Mass lumping generally mitigates but does not completely eliminate outlier frequencies from the spectrum. Thus, it is usually combined with dedicated outlier removal techniques. Unfortunately, the methods described in [14, 22, 16, 12] are only applicable to highly structured problems rarely met in practical applications. Moreover, numerical experiments show that predefined penalization terms barely help remove outliers and must instead be tailored to the specific problem at hand. For this reason, we present in this section a robust and algebraic outlier removal technique based on low-rank perturbations. Our strategy consists in deflating the spectrum from its largest eigenvalues, while preserving the smallest ones. Of course, this choice of scaling assumes that the dynamics are completely resolved by the low-frequency part of the spectrum, which is often (nearly) the case. We first recall some preliminary results, providing the theoretical foundation of the method.
Lemma 4.1 ([52, Theorem VI.1.15]).
@@
where $\operatorname{\operatorname{diag}}(0,f(D_{2})),\operatorname{\operatorname{diag}}(0,g(D_{2}))\in\mathbb{R}^{n\times n}$ are the block diagonal matrices obtained by appending zeros to $f(D_{2})$ and $g(D_{2})$, respectively, and $U=(U_{1},U_{2})$ is the matrix of eigenvectors. Verifying that $\mathbf{u}_{i}$ is an eigenvector of $(\bar{A},\bar{B})$ for $i=1,\dots,n$ is straightforward. Moreover, since the matrix pencils $(\bar{A},\bar{B})$ and $(U^{T}\bar{A}U,U^{T}\bar{B}U)$ are equivalent [53, Chapter 15],
$\Lambda(\bar{A},\bar{B})=\Lambda(U^{T}\bar{A}U,U^{T}\bar{B}U)=\Lambda(D+\operatorname{\operatorname{diag}}(0,f(D_{2})),I+\operatorname{\operatorname{diag}}(0,g(D_{2})))=\{\lambda_{k}\}_{k=1}^{n-r}\cup\left\{\frac{\lambda_{k}+f(\lambda_{k})}{1+g(\lambda_{k})}\right\}_{k=n-r+1}^{n},$
where the second equality follows from Lemma 4.1.
@@
The result, graphically illustrated in Figure 4.1, consists in shaving off the upper part of the spectrum. Note that in our context $\lambda_{n-r}$ is the largest “regular” (or non-outlier) eigenvalue. In principle, it could be replaced with a cutoff value, as suggested in [27, 28], to avoid computing an additional eigenvalue. However, choosing $\lambda_{n-r}$ preserves the eigenvalue numbers and prevents spurious eigenfunctions from moving to the lower to intermediate frequency part of the spectrum. Thus, we prefer computing this additional eigenvalue, which in practice barely introduces any overhead.
@@
On the one hand, Theorem 4.3 circumvents the lack of robustness from predefined perturbations terms. On the other hand, the cost for its implementation is also much higher since it involves a few eigenpairs, which must be computed on a case by case basis. We first focus on the rationale of our method and discuss its computational cost more thoroughly in Section 4.2.
@@
Since $g(D_{2})$ is diagonal (and nonsingular), $(g(D_{2})^{-1}+I_{r})^{-1}$ can be formed explicitly. Thus, solving a linear system with the scaled mass matrix only requires solving a linear system with the lumped mass matrix and computing a matrix-vector multiplication with a low-rank matrix. While the former uses standard techniques, as described in Section 3.5, the latter only requires $O(rn)$ additional flops. Thus, if the perturbation’s rank is relatively small, these matrix-vector multiplications do not introduce any significant overhead. Due to repeated matrix-vector multiplications with the stiffness matrix in the time stepping scheme, the cost incurred by instead scaling the stiffness matrix is the same and eventually the choice is just a matter of taste. The method described herein is very general and could even be beneficial for practitioners using the consistent mass. Indeed, the scaling does not effect the smallest eigenpairs and therefore preserves the higher order convergence characterizing the consistent mass.

### v2

@@ 3.2 Block mass lumping
In general, for nontrivial geometries and coefficients, the mass matrix cannot be expressed as a Kronecker product and the techniques described in [22, 16, 12] are not straightforwardly applicable. In [32], the authors suggested computing the nearest Kronecker product approximation and substituting it to the consistent mass. However, the loss of accuracy induced by the approximation depends on the singular value decay of a reordered matrix and is independent of the discretization parameters. Not only may the approximation be rather crude but combining it with mass lumping also does not give a theoretical guarantee of improving the CFL, although it was observed in all practically relevant cases. The primary objective of this section is to generalize the mass lumping techniques in [32] to complex geometries and address the aforementioned issues. These techniques guarantee an improvement of the CFL condition and their implementation is non-intrusive. Although they do not improve the convergence rate, they might still significantly improve the accuracy and remain relevant, especially at a time when most practical simulations employ quadratic or cubic discretizations. The key observation for achieving these results is that, although the mass matrix generally cannot be expressed as a Kronecker product, it still inherits some favorable structure from the tensor product basis functions which we will exploit for designing algebraic mass lumping techniques. We will first describe this structure in detail for the single-patch case and then propose a block generalization of the methods presented in [32]. From there, the multipatch case naturally follows.
@@
This hierarchical notion of bandedness is related to the standard notion of bandedness through the bandwidths and block sizes. In the rest of the paper we will denote $r_{k}=\prod_{j=k+1}^{d}n_{j}$ the size of the blocks on the $k$th hierarchical level (with $k<d$). On the finest level, the “blocks” reduce to scalars and we set $r_{d}=1$.
@@
Definitions 3.4 and 3.5 only provide a description of the sparsity. In order to build a working framework, they must be complemented with Definitions 3.7 and 3.8 that describe both symmetry and spectral properties.
@@
$\mathcal{S}_{(n_{1},n_{2},\dots,n_{d})}=\{\mathcal{B}\in\mathcal{S}_{n}\colon B_{i,j}\in\mathcal{S}_{(n_{2},\dots,n_{d})}\},$
$\mathcal{S}_{(n_{1},n_{2},\dots,n_{d})}^{+}=\{\mathcal{B}\in\mathcal{S}_{n}^{+}\colon B_{i,j}\in\mathcal{S}_{(n_{2},\dots,n_{d})}\},$
@@
$\mathcal{S}_{(n_{1},n_{2},\dots,n_{d})}=\mathcal{S}_{(n_{1},n_{2},\dots,n_{d-1},r_{d-1})}\subseteq\mathcal{S}_{(n_{1},n_{2},\dots,n_{d-2},r_{d-2})}\subseteq\dots\subseteq\mathcal{S}_{(n_{1},r_{1})}\subseteq\mathcal{S}_{n},$
$\mathcal{S}_{(n_{1},n_{2},\dots,n_{d})}^{+}=\mathcal{S}_{(n_{1},n_{2},\dots,n_{d-1},r_{d-1})}^{+}\subseteq\mathcal{S}_{(n_{1},n_{2},\dots,n_{d-2},r_{d-2})}^{+}\subseteq\dots\subseteq\mathcal{S}_{(n_{1},r_{1})}^{+}\subseteq\mathcal{S}_{n}^{+}.$
@@
We prove the statements from left to right. Let $\mathcal{B}\in\mathcal{S}_{(n_{1},n_{2},\dots,n_{d})}$. By definition, on the $(d-1)$th hierarchical level the matrices are in $\mathcal{S}_{n_{d}}=\mathcal{S}_{r_{d-1}}$ and the first equality trivially follows. On the $(d-2)$th level the matrices are in $\mathcal{S}_{(n_{d-1},r_{d-1})}\subseteq\mathcal{S}_{r_{d-2}}$. Thus, $\mathcal{B}\in\mathcal{S}_{(n_{1},n_{2},\dots,n_{d-2},r_{d-2})}$. We then repeatedly apply the same argument by moving up the hierarchical structure and realizing that at level $k$ (with $20\leq k\leq d-$ 2) the matrices are in $\mathcal{S}_{(n_{k+1},r_{k+1})}\subseteq\mathcal{S}_{r_{k}}$. The proof of the second statement is completely analogous. ∎
@@
$\mathbf{x}^{T}(\mathcal{L}(\mathcal{B})-\mathcal{B})\mathbf{x}$
$=\sum_{i=1}^{n_{1}}\mathbf{x}_{i}^{T}\left(\sum_{j=1}^{n_{1}}B_{i,j}\right)\mathbf{x}_{i}-\sum_{i,j=1}^{n_{1}}\mathbf{x}_{i}^{T}B_{i,j}\mathbf{x}_{j}$
@@
$\lambda_{k}(\mathcal{A},\mathcal{P}_{1})\leq\lambda_{k}(\mathcal{A},\mathcal{P}_{2})\leq\dots\leq\lambda_{k}(\mathcal{A},\mathcal{P}_{n_{1}-1})\leq\lambda_{k}(\mathcal{A},\mathcal{P}_{n_{1}})\quad 1\leq k\leq n.$
Clearly, block diagonal matrices such as $\mathcal{P}_{1}$ are very appealing given that linear systems can be solved in parallel for each block. The block tridiagonal case can still be treated efficiently by a block forward elimination and backward substitution algorithm; see e.g. [49, Section 4.5.1] for the details.
@@ 3.3 Hierarchical mass lumping
Figure 3.2, for example, shows the sparsity pattern of a matrix $\mathcal{B}\in\mathcal{S}_{(n_{1},n_{2},n_{3})}^{+}$ together with its hierarchical lumped mass matrices $\mathcal{H}_{k}$ for $k=1,2,3$. By construction, hierarchical mass lumping reduces the bandwidth down the hierarchical structure: if $\mathcal{B}$ is $d$-level banded with bandwidths $(b_{1},b_{2},\dots,b_{d})$, then $\mathcal{H}_{k}$ is $d$-level banded with bandwidths $(0,\dots,0,b_{k+1},\dots,b_{d})$. Note that the numbering of $\mathcal{H}_{k}$ refers to the hierarchical level and hence typically ranges from $1$ to $d$ (contrary to the numbering of $\mathcal{P}_{i}$, which ranges from $1$ to $n_{1}$). Similarly to (3.5), hierarchical lumped matrices also satisfy an order relation.
@@ 3.4 Multipatch mass lumping
First recall that for any symmetric matrices $A,B\in\mathbb{R}^{n\times n}$ and any $V\in\mathbb{R}^{n\times m}$, if $A\succeq B$, then $V^{T}AV\succeq V^{T}BV$ [50, Theorem 7.7.2(a)]. The result then immediately follows since for any $1\leq r\leq N_{p}$ and any index $1\leq i<n_{1}$,
@@
The statement then follows from repeated application of the previous relation. ∎
@@
The purpose of mass lumping is first and foremost to reduce the block bandedness of the mass matrix and guarantee a CFL condition that cannot be worse than the original one. However, generally speaking, mass lumping does not significantly improve the CFL while it might undermine the accuracy. Multipatch problems are not exempt and the issue already originates on the single-patch level, before local matrices are merged into the global one. This merging process is identical to the assembly procedure of classical finite element methods, which is not surprising given the analogy between patches and elements. Thus, the proof of the following lemma is analogous to the standard finite element case (see e.g. [51, 52, 53]).
@@
The upper bound of Lemma 3.23 may be quite tight and is an incentive for acting on the single-patch matrices prior to assembly.
@@ 3.5 Solving linear systems with the lumped mass matrix
In the single-patch case, linear systems with the lumped mass matrices are conveniently solved using sparse Cholesky factorizations; i.e. $P=LL^{T}$, where $L$ is a lower triangular matrix. It is well-known that the bandwidth of the Cholesky factor $L$ cannot be larger than the bandwidth of $P$. For this reason, many techniques for minimizing the fill-in actually minimize the bandwidth of a permuted matrix. The mass lumping techniques discussed in this work reduce the bandwidth of the consistent mass, thereby significantly accelerating sparse direct solvers. Indeed, a direct application of Lemma 3.6 shows that the bandwidth of $\mathcal{H}_{k}$ is
@@
where $S=X-C^{T}D^{-1}C$ is the Schur complement of $D$ in $\mathcal{P}$ and $\tilde{\mathbf{g}}=\mathbf{g}-C^{T}D^{-1}\mathbf{f}$. Contrary to the consistent mass matrix (which features a similar structure), the Schur complement of the lumped mass matrix can be formed explicitly and cheaply owing to its sparsity and simple block diagonal structure. Once the Schur complement is formed, which is done once and for all, (3.6) can be solved by backward substitution. This strategy is advantageous given that within time stepping schemes, one must solve a sequence of linear systems and not just a single one.
@@ 4.1 Deflation techniques
Mass lumping generally mitigates but does not completely eliminate outlier frequencies from the spectrum. Thus, it is usually combined with dedicated outlier removal techniques. Unfortunately, the methods described in [14, 22, 16, 12] are only applicable to highly structured problems rarely met in practical applications. Moreover, numerical experiments show that predefined penalization terms barely help remove outliers and must instead be tailored to the specific problem at hand. For this reason, we present in this section an algebraic outlier removal technique based on low-rank perturbations that is robust with respect to the geometry. Our strategy consists in deflating the spectrum from its largest eigenvalues, while preserving the smallest ones. Of course, this choice of scaling assumes that the dynamics are completely resolved by the low-frequency part of the spectrum, which is often (nearly) the case. We first recall some preliminary results, providing the theoretical foundation of the method.
Lemma 4.1 ([55, Theorem VI.1.15]).
@@
where $\operatorname{\operatorname{diag}}(0,f(D_{2})),\operatorname{\operatorname{diag}}(0,g(D_{2}))\in\mathbb{R}^{n\times n}$ are the block diagonal matrices obtained by appending zeros to $f(D_{2})$ and $g(D_{2})$, respectively, and $U$ is the matrix of eigenvectors. Verifying that $\mathbf{u}_{i}$ is an eigenvector of $(\bar{A},\bar{B})$ for $i=1,\dots,n$ is straightforward. Moreover, since the matrix pencils $(\bar{A},\bar{B})$ and $(U^{T}\bar{A}U,U^{T}\bar{B}U)$ are equivalent [56, Chapter 15],
$\Lambda(\bar{A},\bar{B})$ $=\Lambda(U^{T}\bar{A}U,U^{T}\bar{B}U)$
$=\Lambda(D+\operatorname{\operatorname{diag}}(0,f(D_{2})),I+\operatorname{\operatorname{diag}}(0,g(D_{2})))$
=$=\{\lambda_{k}\}_{k=1}^{n-r}\cup\left\{\frac{\lambda_{k}+f(\lambda_{k})}{1+g(\lambda_{k})}\right\}_{k=n-r+1}^{n},$
where the second equality follows from Lemma 4.1. ∎
@@
The result, graphically illustrated in Figure 4.1, consists in shaving off the upper part of the spectrum. Note that in our context $\lambda_{n-r}$ is the largest “regular” (or non-outlier) eigenvalue. In principle, it could be replaced with a cutoff value, as suggested in [27, 28], to avoid computing an additional eigenvalue. However, choosing $\lambda_{n-r}$ preserves the eigenvalue numbers and prevents spurious eigenfunctions from moving to the lower to intermediate frequency range. Thus, we prefer computing this additional eigenvalue, which in practice barely introduces any overhead.
@@
On the one hand, Theorem 4.3 circumvents the lack of robustness from predefined perturbations terms. On the other hand, the cost for its implementation is also much higher since it involves a few eigenpairs, which must be computed on a case-by-case basis. We first focus on the rationale of our method and discuss its computational cost more thoroughly in Section 4.2.
@@
Since $g(D_{2})$ is diagonal (and nonsingular), $(g(D_{2})^{-1}+I_{r})^{-1}$ can be formed explicitly. Thus, solving a linear system with the scaled mass matrix only requires solving a linear system with the lumped mass matrix and computing a matrix-vector multiplication with a low-rank matrix. While the former uses standard techniques, as described in Section 3.5, the latter only requires $O(rn)$ additional flops. Thus, if the perturbation’s rank is relatively small, these matrix-vector multiplications do not introduce any significant overhead. Due to repeated matrix-vector multiplications with the stiffness matrix in the time stepping scheme, the cost incurred by instead scaling the stiffness matrix is the same and eventually the choice is just a matter of taste. Although stability concerns have been raised for the Woodbury identity [60], instabilities were never experienced during our tests. The method described herein is very general and could even be beneficial for practitioners using the consistent mass. Indeed, the scaling does not affect the smallest eigenpairs and therefore preserves the higher order convergence characterizing the consistent mass. However, the eigenpairs must then be computed to high precision, which is often not necessary when using a lumped mass whose small frequencies are only second order accurate.